51 = 400e^(−0.6x) I want to know how to solve for x using logarithms.
Thank you for your time!!!

Answer by lwsshak3(6472)

51 = 400e^(−0.6x)
51/400=e^(-0.6x)
take ln of both sides
ln 51-ln 400=-0.6xlne
log of base=1
-2.0596=-0.6x
x=-2.0596/-0.6
x=3.4327
Check:
400e^(−0.6x=400e^(−0.6*3.4327)=400*1.275=51.00